A student applies for two different scholarships. The probability of receiving the first scholarship is 0.3 and the probability of receiving the second is 0.4. What is the probability of the student receiving one scholarship? The decisions are made independently.
Answer by stanbon(75887)

The probability of receiving the first scholarship is 0.3 and the probability of receiving the second is 0.4. What is the probability of the student receiving one scholarship? The decisions are made independently.
--------------------
Prob(receives both) = 0.3*0.4 = 0.12
Prob(receives none) = 0.7*0.6 = 0.42
Prob(receives one) = 1 -0.12-0.42 = 1 = 0.54 = 0.46
-------------------
Cheers,
Stan H.